# ClaimVer: Explainable Claim-Level Verification and Evidence Attribution of Text Through Knowledge Graphs

### v1

@@ Abstract
In the midst of widespread misinformation and disinformation through social media and the proliferation of AI-generated texts, it has become increasingly difficult for people to validate and trust information they encounter. Many fact-checking approaches and tools have been developed, but they often lack appropriate explainability or granularity to be useful in various contexts. A text validation method that is easy to use, accessible, and can perform fine-grained evidence attribution has become crucial. More importantly, building user trust in such a method requires presenting the rationale behind each prediction, as research shows this significantly influences people’s belief in automated systems. It is also paramount to localize and bring users’ attention to the specific problematic content, instead of providing simple blanket labels. In this paper, we present ClaimVer, a human-centric framework tailored to meet users’ informational and verification needs by generating rich annotations and thereby reducing cognitive load. Designed to deliver comprehensive evaluations of texts, it highlights each claim, verifies it against a trusted knowledge graph (KG), presents the evidence, and provides succinct, clear explanations for each claim prediction. Finally, our framework introduces an attribution score, enhancing applicability across a wide range of downstream tasks.
@@ 1 Introduction
Misinformation and disinformation are longstanding issues, but the proliferation of AI tools that can generate information on demand has amplified these issues. Tools for fact-checking are not keeping pace with sophisticated text generation techniques. Even when they are effective, they lack appropriate explainability and granularity to be useful to users. Studies have shown that explanations are crucial for users to build trust in AI systems Rechkemmer and Yin (2022); Weitz et al. (2019); Shin (2021). There is a need for a novel human-centric approach to text verification that provides usable and appropriately granular explanations that can not only inform but also educate the user.
Most fact-checkers, including widely used ones in deployment, issue blanket predictions that can lead to user misunderstanding. For instance, in Figure 1 (A), we observe that HealthFeedback, [^2] https://healthfeedback.org/, a fact-checker for medical text, indicates that a misleading statement about the increase in Autism is inaccurate. However, there are multiple claims made in that text, which are not addressed by this tool. In fact, research does show that Autism cases have increased, but this is mostly attributed to increased testing Russell et al. (2015). Our method accurately breaks down the text into multiple claims and shows that the specific claim that vaccines are causing autism is indeed incorrect, attributing it to a fact from the Wikidata Vrandečić and Krötzsch (2014). It also provides a clear rationale as to why the first two claims cannot be determined, as there’s no conclusive evidence present in the KG. Such granular predictions, supported by justifications, significantly improve user confidence Rechkemmer and Yin (2022); Weitz et al. (2019); Shin (2021).
@@
The assumption of one-to-one mapping between input and reference texts, prevalent in previous methods Rashkin et al. (2023); Yue et al. (2023); Thorne et al. (2019); Aly et al. (2021), does not hold if the given text consists of claims that can be mapped to more than one source. In contrast, utilizing a KG, which represents a consolidated body of knowledge, results in a more comprehensive evaluation. While most previous methods may not support scenarios with information spread across various references, querying a KG can yield triplets originally sourced from multiple documents. Additionally, procuring the specific spans of text required to evaluate claims, from large text sources that may span several pages, presents many challenges. On the other hand, a KG captures only the most important relationships as nodes and links, and offers a more efficient way to evaluate the claims.
Prior methods that depend on document indices or vector databases are not easy to maintain or audit. In contrast, existing trusted KGs that are constructed through human curation provide an effective and human-centered approach for evaluating text at scale. Therefore, we leverage KGs to build a framework that realizes our goal of performing fine-grained text verification and evidence attribution. Our framework also generates insights that boost user awareness, thereby fostering increased trust in automated systems.
@@ 2 Related Work
Our method integrates fact verification and evidence attribution; therefore, we discuss recent advancements in both domains in this section.
@@ 2.1 Fact Verification
In Lee et al. (2020), the authors show that the inherent knowledge of LLMs could be used to perform fact verification. Other works Yao et al. (2022); Jiang et al. (2023b) have shown that using external knowledge is helpful for many reasoning-intensive tasks, and report enhanced performance on HotPotQA Yang et al. (2018) and FEVER Thorne et al. (2018). A wide variety of studies have established LLMs are suitable for fact verification. For example, Dong and Smith (2021) enhanced accuracy of table-based fact verification by incorporating column-level cell rank information into pre-training. In FactScore, authors Min et al. (2023), introduce a new evaluation that breaks a long-form text generated by large language models (LMs) into individual atomic facts and calculates the proportion of these atomic facts that are substantiated by a credible knowledge base.
@@ 2.2 Evidence Attribution
The distinction between evidence attribution and fact verification lies in the emphasis on identifying a source that can be attributed to the information. This task is becoming increasingly important, as generative models produce useful and impressive outputs, but without a frame of reference to validate them. In Rashkin et al. (2023), the authors present a framework named AIS (Attributable to Identified Sources) that specifies annotation guidelines and underlines the importance of attributing text to an external, verifiable, and independent source. Yue et al. (2023) demonstrate that LLMs can be utilized for automatic evaluation of attribution, operationalizing the guidelines presented in Rashkin et al. (2023). However, both of these works are primarily designed for the question-answering (QA) task, a primary end-user application for LLMs like ChatGPT Achiam et al. (2023). In contrast, our method is not restricted to QA and is designed to work with text in general. Furthermore, while these previous studies focus on sentence or paragraph levels, our approach extends to a more detailed and granular level of analysis.
@@ 3.3 Objective Function
  • Attributable: The triplets fully support the claim.
  • Extrapolatory: The triplets lack sufficient information to evaluate the claim.
  • Contradictory: The triplets contradict the claim.
@@
  • $input\_text$: The input text containing claim(s).
  • $ret\_triplets$: A set of retrieved triplets for the input text.
  • $claim\_span_{i}$: The $i^{th}$ claim extracted as a substring from $input\_text$.
  • $claim\_pred_{i}$: The predicted label for $claim\_span_{i}$.
  • $rel\_triplets_{i}$: A relevant subset of $ret\_triplets$ supporting, refuting, or are extrapolatory for $claim\_span_{i}$.
@@
  • $n$: The total number of claims automatically extracted from $input\_text$.
@@ 3.4 Fine-tuning LLMs
In order to perform attribution at the claim level, we need to fine-tune LLMs specifically for the proposed objective function (see §3.3) using a custom dataset. This is necessary because, as of this writing, even the state-of-the-art model, OpenAI’s GPT-4 Achiam et al. (2023), does not perform satisfactorily right out of the box. Further details on the dataset are provided in §4. On every prediction, a membership check for relevant triplets is performed for additional verification.
We selected five open-source LLMs with diverse sizes, ranging from 2.7B parameters to 13B parameters, to perform the fine-tuning: Phi-2 2.7B Javaheripi et al. (2023), Mistral-Instruct 7B Jiang et al. (2023a), Zephyr-Beta 7B Tunstall et al. (2023), Solar-Instruct 10.7B Kim et al. (2023), and Llama2-chat 13B Touvron et al. (2023). The models were fine-tuned using LoRA Hu et al. (2021) with 4-bit quantization and adapters with rank 8 Dettmers et al. (2024). The context length was set to 1024 tokens. All models converged after 2 epochs, and high ROUGE-L Lin (2004) scores greater than 0.635 were achieved for each model. The instruction prompt used for fine-tuning is presented in Figure. 3.
@@
Figure 3: Instruction prompt for operationalizing objective function.
@@ 3.5.1 Claim Scores
$\text{cs}(y_{i})=\begin{cases}2&\text{if }y_{i}=\text{Attributable}\\ 1&\text{if }y_{i}=\text{Extrapolatory}\text{ and }|\textit{rel\_triplets}_{i}|>0\\ 0&\text{if }y_{i}=\text{Extrapolatory}\text{ and }|\textit{rel\_triplets}_{i}|=0\\ 0&\text{if }y_{i}=\text{No attribution}\\ -1&\text{if }y_{i}=\text{Contradictory}\end{cases}$ (2)
@@
For each claim, we assign a score that reflects the level of its validity, ranging from -1 (contradictory) to 2 (attributable). If a claim is predicted to be extrapolatory, yet has one or more relevant triplets, we assign that claim a score of 1, as there is still relevant information available even though it may not be sufficient to completely support or refute. However, if there are no triplets at all, along with an extrapolatory prediction, we assign 0 as it does not add much information. While decomposing claims, the model might occasionally omit words, typically stop-words, and we assign 0 in those cases as well.
@@ 3.5.2 Triplets Match Score (TMS)
This score reflects the extent of the match between the relevant triplets and the corresponding claim, and it can also serve as a proxy for prediction confidence. Even though the prediction is made at the claim level, the triplets match score considers word-level matches in the computation. It can be computed as follows:
$TMS($ $E(claim\_span_{i}),E(rel\_triplet_{i}))=$ (3)
$\alpha\cdot SS(E(claim\_span_{i}),E(rel\_triplet_{i}))$
$+\beta\cdot EPR(E(claim\_span_{i}),E(rel\_triplet_{i}))$
@@ 4 Dataset
Table 2: Examples of claim-level attribution by the proposed method. The first column shows the numbered claims in the input text. Second column lists relevant triplets for each claim. Predictions and Triplets Match Score (TMS) are in the third column, while the rationale behind each prediction is in the fourth column. The Knowledge Graph Attribution Score (KAS) is shown in the last column. Model: Solar-Instruct.
@@
Open-domain Question Answering (QA) datasets, such as WikiQA Yang et al. (2015), HotPotQA Yang et al. (2018), PopQA Mallen et al. (2022), and EntityQuestions Sciavolino et al. (2021), as well as Fact Verification datasets like FEVER Thorne et al. (2019), FEVEROUS Aly et al. (2021), TabFacT Chen et al. (2019), and SEM-TAB-FACTS Wang et al. (2021), provide texts along with corresponding reference contexts or attributable information sources. However, these datasets significantly differ from the type of data required to train and test our proposed objective function, primarily due to two major factors: First, these datasets predominantly offer samples that are inherently attributable. To address this limitation, prior work Yue et al. (2023) in attribution evaluation introduced new samples by modifying correct answers to generate contradictory instances. However, this adjustment alone is not adequate for our use case because, our method requires attribution at the claim level, and necessitates the automatic decomposition of claims. Consequently, as this task represents a novel challenge, we developed a new dataset that enables effective training and testing of the objective function.
@@
Considering the choice of our KG, which is Wikidata Vrandečić and Krötzsch (2014), we opted for WikiQA Yang et al. (2015) as it is closely associated with the Wiki ecosystem. Given that our method is designed for text validation in general, not limited to question answering, we retain only answers and discard the questions. Subsequently, we processed the answers following the steps detailed in Section 3.1, selecting entries containing two or more Wiki entities. This approach resulted in the exclusion of most single-word answers and other responses that are dependent on their corresponding questions and may lack comprehensibility without them.
@@
We utilize GPT-4 Achiam et al. (2023) to generate the initial version of the ground truth, as knowledge distillation Gou et al. (2021) has proven to be an effective strategy. Although GPT-4 can adhere to the instructions (refer to Figure 3) to a reasonable degree and responds in the required format with all necessary keys, it still underperforms in the overall task. The most frequent issue observed is the erroneous assignment of prediction labels. After post-processing, we conducted manual checks to ensure only high-quality samples were retained, as research indicates that high alignment can be achieved with as few as 1,000 samples, provided they are of superior quality Zhou et al. (2023).
@@
The final dataset is comprised of two splits: the training split, based on the training split of WikiQA Yang et al. (2015), and a test split, derived from both the test and validation splits. The training split contains 3,400 samples, and since some entries feature multiple claims, there are a total of 5,343 claims within this split. Similarly, the test split includes 1,000 samples and 1,675 claims. The label counts for the claims are tabulated in Table 1.
@@ 5 Experiments and Results
In this section, we present the evaluation of our claim-level attribution method. The performance metrics of the fine-tuned LLMs, which operationalize the objective function, are presented in Tables 3 and 4. In Table 3, we observe that all models converge and achieve sufficiently high ROUGE-L and ROUGE-1 scores, with Solar-Instruct achieving the highest of 0.655 and 0.693 respectively. We also observe that the smaller model, Phi2 with just 2.7B parameters, is also sufficiently compatible for this task as it attained a decent ROUGE-L score of 0.635.
The average accuracies on the test set, however, vary significantly across the models, with Solar-Instruct reporting the highest at 89.31%. The reason behind this variation is that these scores account for both sub-tasks of the objective function (refer §3.3): decomposing input text into claims and generating predictions for each claim. The first task, decomposing text into multiple claims, is somewhat subjective, and there could be multiple valid approaches due to linguistic complexities. We impose a strict strategy while computing accuracy: the text span of the claim, the identified relevant triplets, and the prediction label must all exactly match the ground truth to be considered accurate. Since these LLMs may decompose the claims slightly differently, as multiple valid options are possible, the accuracy values may appear low even while the objective function is correctly executed. For instance, example 4 in Table 2 has been decomposed into three claims, but the first could arguably be further decomposed to verify whether Blagnac is in France, and whether it is a suburb of Toulouse. Controlling the precise manner of decomposition is challenging, and might necessitate an additional step before the prediction step, involving separate processing for each claim. However, this option could prove to be impractical, as the number of LLM queries could increase exponentially. While the overall accuracy may not fully reflect the models’ performance due to the combined assessment of the two sub-tasks, focusing solely on the prediction task could offer better insights into how the models are performing in terms of categorization.
In Table 4, the second column indicates number of claims with text spans exactly matching the ground truth responses. Columns 3 to 6 present the accuracy, precision, recall, and F1 scores for these matching claims. The most performant model is Solar-Instruct, with 1052 exact matches out of 1675 claims in the test set. Across all models, the classification scores in all metrics are above 98%, which clearly demonstrates that the models can reliably differentiate between the classes attributable, extrapolatory, and contradictory.
@@
Model Size ROUGE-L ROUGE-1 Avg. Acc.
Phi-2 2.7B 0.635 0.673 75.16%
Mistral-Instruct 7B 0.645 0.680 83.04%
Zephyr-Beta 7B 0.638 0.676 79.88%
Solar-Instruct 10.7B 0.655 0.693 89.31%
Llama2-Chat 13B 0.6395 0.677 79.41%
Table 3: ROUGE scores and average accuracies on the test set ($n=1,000$).
@@
Phi-2 819 98.29 98.33 98.29 98.26
Mistral-Instruct 928 99.35 99.36 99.35 99.35
Zephyr-Beta 757 98.34 98.38 98.34 98.32
Solar-Instruct 1052 99.80 99.81 99.80 99.80
Llama2-Chat 869 99.53 99.54 99.53 99.53
Table 4: Scores on matching claims in the test set ($n=1675$). #MC: number of matching claims.

### v2

@@ Abstract
In the midst of widespread misinformation and disinformation through social media and the proliferation of AI-generated texts, it has become increasingly difficult for people to validate and trust information they encounter. Many fact-checking approaches and tools have been developed, but they often lack appropriate explainability or granularity to be useful in various contexts. A text validation method that is easy to use, accessible, and can perform fine-grained evidence attribution has become crucial. More importantly, building user trust in such a method requires presenting the rationale behind each prediction, as research shows this significantly influences people’s belief in automated systems. Localizing and bringing users’ attention to the specific problematic content is also paramount, instead of providing simple blanket labels. In this paper, we present ClaimVer, a human-centric framework tailored to meet users’ informational and verification needs by generating rich annotations and thereby reducing cognitive load. Designed to deliver comprehensive evaluations of texts, it highlights each claim, verifies it against a trusted knowledge graph (KG), presents the evidence, and provides succinct, clear explanations for each claim prediction. Finally, our framework introduces an attribution score, enhancing applicability across a wide range of downstream tasks.
ClaimVer: Explainable Claim-Level Verification and Evidence Attribution of Text Through Knowledge Graphs
Preetam Prabhu Srikar Dammu1, Himanshu Naidu1, Mouly Dewan 1, YoungMin Kim 1, Tanya Roosta2,4,thanks: Work does not relate to position at Amazon., Aman Chadha3,4,11footnotemark: 1, Chirag Shah1 1University of Washington     2UC Berkeley    3Stanford University     4Amazon GenAI
@@ 1 Introduction
Misinformation and disinformation are longstanding issues, but the proliferation of AI tools that can generate information on demand has amplified these issues. Tools for fact-checking are not keeping pace with sophisticated text generation techniques. Even when they are effective, they lack appropriate explainability and granularity to be useful to users. Studies have shown that explanations are crucial for users to build trust in AI systems Rechkemmer and Yin (2022); Weitz et al. (2019); Shin (2021). Therefore, there is a need for a novel human-centric approach to text verification that offers usable and sufficiently granular explanations to inform and educate the user.
Most fact-checkers, including widely used ones in deployment, issue blanket predictions that can lead to user misunderstandings. For instance, in Figure 1 (A), we observe that HealthFeedback [^1] https://healthfeedback.org/, a fact-checker for medical text, indicates that a misleading statement about the increase in Autism is inaccurate. However, there are multiple claims made in that text, which are not addressed by this tool. In fact, research does show that Autism cases have increased, but this is mostly attributed to increased testing Russell et al. (2015). Our method accurately breaks down the text into multiple claims and shows that the specific claim that vaccines are causing autism is indeed incorrect, attributing it to a fact from the Wikidata Vrandečić and Krötzsch (2014). It also provides a clear rationale as to why the first two claims cannot be determined, as there’s no conclusive evidence present in the KG. Such granular predictions, supported by justifications, significantly improve user confidence Rechkemmer and Yin (2022); Weitz et al. (2019); Shin (2021).
@@
The assumption of one-to-one mapping between input and reference texts, prevalent in previous methods Rashkin et al. (2023); Yue et al. (2023); Thorne et al. (2019); Aly et al. (2021), does not hold if the given text consists of claims that can be mapped to more than one source. In contrast, utilizing a KG, which represents a consolidated body of knowledge, results in a more comprehensive evaluation. While most previous methods may not support scenarios with information spread across various references, querying a KG can yield triplets originally sourced from multiple documents. Additionally, procuring the specific spans of text required to evaluate claims from large text sources that may span several pages presents many challenges. In contrast, a KG captures only the most important relationships as nodes and links, providing a more efficient way to evaluate the claims.
@@ 2 Related Work
Our method integrates fact verification and evidence attribution. In this section, we discuss recent advancements in both domains.
@@ 2.1 Fact Verification
In Lee et al. (2020), the authors show that the inherent knowledge of LLMs could be used to perform fact verification. Other works Yao et al. (2022); Jiang et al. (2023b) have shown that using external knowledge is helpful for many reasoning-intensive tasks, and report enhanced performance on HotPotQA Yang et al. (2018) and FEVER Thorne et al. (2018). A wide variety of studies have established that LLMs are suitable for fact verification. For example, Dong and Smith (2021) enhanced accuracy of table-based fact verification by incorporating column-level cell rank information into pre-training. In FactScore, authors Min et al. (2023), introduce a new evaluation that breaks a long-form text generated by large language models (LMs) into individual atomic facts and calculates the proportion of these atomic facts that are substantiated by a credible knowledge base.
@@ 2.2 Evidence Attribution
The distinction between evidence attribution and fact verification lies in the emphasis on identifying a source that can be attributed to the information. This task is becoming increasingly important, as generative models produce useful and impressive outputs, but without a frame of reference to validate them. In Rashkin et al. (2023), the authors present a framework named AIS (Attributable to Identified Sources) that specifies annotation guidelines and underlines the importance of attributing text to an external, verifiable, and independent source. Yue et al. (2023) demonstrate that LLMs can be utilized for automatic evaluation of attribution, operationalizing the guidelines presented in Rashkin et al. (2023). However, both of these works are primarily designed for the question-answering (QA) task. In contrast, our method is not restricted to QA and is designed to work with text in general. Furthermore, while these previous studies focus on sentence or paragraph levels, our approach extends to a more detailed and granular level of analysis.
@@ 3.3 Objective Function
  • Attributable: Triplets fully support the claim.
  • Extrapolatory: Triplets lack sufficient information to evaluate the claim.
  • Contradictory: Triplets contradict the claim.
@@
  • $input\_text$: input text containing claim(s).
  • $ret\_triplets$: retrieved triplets for the input text.
  • $claim\_span_{i}$: $i^{th}$ claim extracted as a substring from $input\_text$.
  • $claim\_pred_{i}$: label predicted for $claim\_span_{i}$.
  • $rel\_triplets_{i}$: relevant subset of $ret\_triplets$ for $claim\_span_{i}$.
@@
  • $n$: total number of claims in $input\_text$.
@@ 3.4 Fine-tuning LLMs
In order to perform attribution at the claim level, we need to fine-tune LLMs specifically for the proposed objective function (see §3.3) using a custom dataset. This is necessary because, as of this writing, even the state-of-the-art model, OpenAI’s GPT-4 Achiam et al. (2023), does not perform satisfactorily right out of the box. Our custom dataset, built using two sequential complex prompts with GPT-4, enables us to fine-tune significantly smaller models. This approach distills the performance of a large proprietary model using a multi-query prompt pipeline into small open-source models with a compact zero-shot prompt. We make the weights of the fine-tuned models publicly available333weights available on HuggingFace.
We selected eight open-source LLMs with diverse sizes, ranging from 2B parameters to 10B parameters, to perform the fine-tuning: Gemma-2B-IT-Chat Team et al. (2024), Phi-3-mini-4k-Chat Javaheripi et al. (2023), Zephyr-7B-Beta-Chat Tunstall et al. (2023), Mistral-7B-v0.3-Chat Jiang et al. (2023a), Llama3-8B-Chat Touvron et al. (2023), and Solar-10.7B-Chat Kim et al. (2023). The models were fine-tuned using LoRA Hu et al. (2021) with 4-bit quantization and adapters with rank 8 Dettmers et al. (2024). The context length was set to 4096 tokens (for additional training details, refer §A.1) All models converged after 2 epochs, and high ROUGE-L Lin (2004) scores greater than 0.658 were achieved for each model. The instruction prompt used for fine-tuning is presented in Figure 3.
@@
Figure 3: Instruction prompt for fine-tuned LLMs.
@@ 3.5.1 Claim Scores
$\text{cs}(y_{i})=\begin{cases}2&\text{if }y_{i}=\text{Attributable}\\ 1&\text{if }y_{i}=\text{Extrapolatory}\text{ and }|\textit{rel\_triplets}_{i}|>0\\ 0&\text{if }y_{i}=\text{Extrapolatory}\text{ and }|\textit{rel\_triplets}_{i}|=0\\ 0&\text{if }y_{i}=\text{No attribution}\\ -1&\text{if }y_{i}=\text{Contradictory}\end{cases}$
@@
For each claim, we assign a score that reflects the level of its validity, ranging from -1 (contradictory) to 2 (attributable). If a claim is predicted to be extrapolatory, yet has one or more relevant triplets, we assign that claim a score of 1, as there is still relevant information available even though it may not be sufficient to completely support or refute the claim. However, if there are no triplets at all, along with an extrapolatory prediction, we assign 0 as it does not add any useful information. While decomposing claims, the model might occasionally omit words, typically stop-words, and we assign 0 in those cases as well.
@@ 3.5.2 Triplets Match Score (TMS)
This score reflects the extent of the match between the relevant triplets and the corresponding claim, and it can also serve as a proxy for the prediction confidence. Even though the prediction is made at the claim level, the triplets match score considers word-level matches in the computation. It can be computed as follows:
[…]
@@ 4 Dataset
Table 2: Examples of claim-level attribution by the proposed method. The first column shows the numbered claims in the input text. Second column lists relevant triplets for each claim. Predictions and Triplets Match Score (TMS) are in the third column, while the rationale behind each prediction is in the fourth column. The Knowledge Graph Attribution Score (KAS) is shown in the last column. Model: Solar-10.7B-Chat.
@@
Open-domain Question Answering (QA) datasets, such as WikiQA Yang et al. (2015), HotPotQA Yang et al. (2018), PopQA Mallen et al. (2022), and EntityQuestions Sciavolino et al. (2021), as well as Fact Verification datasets like FEVER Thorne et al. (2019), FEVEROUS Aly et al. (2021), TabFacT Chen et al. (2019), and SEM-TAB-FACTS Wang et al. (2021a), provide texts along with corresponding reference contexts or attributable information sources. However, these datasets significantly differ from the type of data required to train and test our proposed objective function, primarily due to two major factors: (i) these datasets predominantly offer samples that are inherently attributable, and (ii) consist of atomic claims and/or one-to-one mappings between input and reference texts. To address the first limitation, prior work Yue et al. (2023) in attribution evaluation introduced new samples by modifying correct answers to generate contradictory instances. Yet, this adjustment alone is not sufficient for our use case because our method requires attribution at the claim level, and necessitates the automatic decomposition input text to claims. Consequently, as this task represents a novel challenge, we developed a new dataset that enables effective training and testing of the objective function.
@@
Considering the choice of our KG, which is Wikidata Vrandečić and Krötzsch (2014), we opted for WikiQA Yang et al. (2015) as it is closely associated with the Wiki ecosystem. Given that our method is designed for text validation in general, and is not limited to question answering, we retain only answers and discard the questions. Subsequently, we processed the answers following the steps detailed in Section 3.1, selecting entries containing two or more Wiki entities. This approach resulted in the exclusion of most single-word answers and other responses that are dependent on their corresponding questions and may lack comprehensibility without them.
@@
We utilize GPT-4 Achiam et al. (2023) to generate the initial version of the ground truth. Although GPT-4 can adhere to the instructions (refer to Figure 3) to a reasonable degree and responds in the required format with all necessary keys, it still underperforms in the overall task. The most frequent issue observed is the erroneous assignment of prediction labels. To remedy this issue, we designed a detailed prompt tailored for the given task, incorporating techniques such as few-shot, chain-of-thought Kojima et al. (2022), and other strategies OpenAI (2024); Nori et al. (2023) (full prompt in §A Figure 12). We also conducted manual checks to ensure only high-quality samples were retained, as research indicates that high alignment can be achieved with as few as 1,000 samples, provided they are of superior quality Zhou et al. (2023).
@@
The final dataset is comprised of two splits: the training split, based on the training split of WikiQA Yang et al. (2015), and a test split, derived from both the test and validation splits. The training split contains 3,400 samples, and since some entries feature multiple claims, there are a total of 5,342 claims within this split. Similarly, the test split includes 1,000 samples and 1,677 claims. The label counts for the claims are tabulated in Table 1. The dataset is publicly shared to facilitate further research in this direction444dataset available on HuggingFace.
@@ 5 Experiments and Results
In this section, we present the evaluation of our claim-level attribution method. The performance metrics of the fine-tuned LLMs, which operationalize the objective function, are presented in Tables 3 and 4. In Table 3, we observe that all models converge and achieve sufficiently high ROUGE-L and ROUGE-1 scores, with Mistral-7B-v0.3-Chat achieving the highest of 0.694 and 0.719 respectively. We also observe that the smaller model, Gemma-2B-IT-Chat with just 2B parameters, is also sufficiently compatible for this task as it attained a decent ROUGE-L score of 0.667.
The first task of the proposed objective function (refer §3.3), decomposing text into multiple claims, is somewhat subjective, and there could be multiple valid approaches due to linguistic complexities. For instance, example 4 in Table 2 has been decomposed into three claims, but the first could arguably be further decomposed to verify whether Blagnac is in France, and whether it is a suburb of Toulouse. Controlling the precise manner of decomposition is challenging, and might necessitate an additional step before the prediction step, involving separate processing for each claim. However, this option could prove to be impractical, as the number of LLM queries could increase exponentially.
To accurately compute classification performance, we impose a strict strategy: the text span of the claim, the identified relevant triplets, and the prediction label must all exactly match the ground truth to be considered accurate. In Table 4, the second column indicates number of claims with text spans exactly matching the ground truth responses. Columns 3 to 6 present the accuracy, precision, recall, and F1 scores for these matching claims. The most performant model is Solar-10.7B-Chat, with 1031 exact matches out of 1677 claims in the test set. Additionally, the classification scores in all metrics are above 89%, which clearly demonstrates that the model can reliably differentiate between the classes attributable, extrapolatory, and contradictory.
@@
Model Size ROUGE-L ROUGE-1
Gemma-2B-IT-Chat 2B 0.667 0.692
Phi-3-mini-4k-Chat 4B 0.658 0.685
Zephyr-7B-Beta-Chat 7B 0.686 0.712
Vicuna-7B-v1.5-Chat 7B 0.676 0.700
Mistral-7B-v0.3-Chat 7B 0.694 0.719
Gemma-7B-IT-Chat 7B 0.678 0.703
Llama3-8B-Chat 8B 0.679 0.705
Solar-10.7B-Chat 10B 0.689 0.714
Table 3: ROUGE scores on the test set ($n=1,000$).
@@
Gemma-2B-IT-Chat 895 77.09 77.20 77.09 74.24
Phi-3-mini-4k-Chat 882 72.22 78.10 72.22 72.86
Zephyr-7B-Beta-Chat 978 85.89 87.41 85.89 86.16
Vicuna-7B-v1.5-Chat 898 79.62 78.83 79.62 78.84
Mistral-7B-v0.3-Chat 1002 86.63 87.03 86.63 86.73
Gemma-7B-IT-Chat 940 82.87 84.09 82.87 83.17
Llama3-8B-Chat 959 80.92 85.48 80.92 81.36
Solar-10.7B-Chat 1031 89.23 89.52 89.23 89.30
Table 4: Scores on matching claims in the test set ($n=1677$). #MC: number of matching claims.